# Accuracy of Discretely Sampled Stochastic Policies in Continuous-time Reinforcement Learning

### v1

@@ 7. Proofs of Main Results
$\leq C_{p}\mathbb{E}\left[\left(\int_{0}^{t}|\Delta X_{u}|^{2}\mathrm{d}u\right)^{\frac{p}{2}}\right]\leq C_{p}\mathbb{E}\left[\int_{0}^{t}|\Delta X_{u}|^{p}\mathrm{d}u\right].$
Thus, we have
@@
$+C_{p}\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))\mathrm{d}s-\int_{0}^{u}\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\mathrm{d}s\right|^{p}\right].$
We focus on the third term of the above equation. Denote $U_{s}:=b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})$. Under (H.6), we have
@@
Denote $\mathbb{E}_{\xi}\coloneqq\mathbb{E}\left[\cdot|\mathcal{F}^{W}\right]$ and note that $\xi$ is independent of $\mathcal{F}^{W}$.
$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\leq C_{p}\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{\delta(u)}U_{s}\mathrm{d}s\right|^{p}\right]+C_{p}\mathbb{E}\left[\mathbb{E}_{\xi}\left[\sup_{0\leq u\leq t}\left|\int_{\delta(u)}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\right].$ (7.34)
It follows from (7.33) and Lemma 7.3 that
$\mathbb{E}\left[\mathbb{E}_{\xi}\left[\sup_{0\leq u\leq t}\left|\int_{\delta(u)}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\right]$ $\leq C_{p}\mathbb{E}\left[\left(1+\sup_{0\leq t\leq T}\left|\widetilde{X}_{t}\right|^{p}+\mathbb{E}\left[\max_{1\leq j\leq n}\left(d_{A}(\phi(0,0,\xi_{j}),a_{0})\right)^{p}\right]\right)|\mathscr{G}|^{p}\right]$ (7.35)
@@
where in the last inequality, we used the fact that $|\mathscr{G}|^{1/2}\log(1+n)\leq C_{G}$.
As for the first term on the right hand side of (7.34), denote $M_{j}:=\int_{t_{j}}^{t_{j+1}}U_{s}\mathrm{d}s$ and $\mathcal{F}_{i}:=\sigma(\xi_{j}:0\leq j\leq i-1)\lor\mathcal{F}^{W}$. Note that $\left\{(\sum_{j=0}^{i-1}M_{j},\mathcal{F}_{i})\right\}_{i=1}^{n}$ is a martingale. It follows from Doob’s inequality that
$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{\delta(u)}U_{s}\mathrm{d}s\right|^{p}\right]=\mathbb{E}\left[\max_{1\leq i\leq n}\Big{|}\sum_{j=0}^{i-1}M_{j}\Big{|}^{p}\right]\leq C_{p}\mathbb{E}\left[\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}^{p}\right],$ (7.36)
where $C_{p}$ is a constant depend on $p$.
By (H.6) and relation (7.33), $M_{j}$ is sub-exponential conditional on $\mathcal{F}^{W}$ and with a little abuse of notations,
$\|M_{j}\|_{\psi_{1}}\leq C\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\|1\|_{\psi_{1}}+\|d_{A}(\phi(0,0,\xi),a_{0})\|_{\psi_{1}}\right)(t_{j+1}-t_{j}),$ (7.37)
where $\|\cdot\|_{\psi_{1}}$ is the Orlicz norm on $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$. It follows from [25, Theorem 6.21] that
$\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\leq C\left(\mathbb{E}_{\xi}\left[\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}\right]+\|\max_{j}|M_{j}|\|_{\psi_{1}}\right).$ (7.38)
For the first term on the right hand side of the above equation, using the fact that $\mathbb{E}_{\xi}\left[M_{i}^{\top}M_{j}\right]=0,i\neq j$, it follows from relation (7.33) and $\psi(0,0,\xi)$ is square integrable that
$\mathbb{E}_{\xi}\bigg{[}\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}\bigg{]}\leq\sqrt{\mathbb{E}_{\xi}\bigg{[}\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}^{2}\bigg{]}}=\sqrt{\sum_{j=0}^{n-1}\mathbb{E}_{\xi}\left[|M_{j}|^{2}\right]}\leq C(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|)|\mathscr{G}|^{\frac{1}{2}}.$ (7.39)
As for the second term, it follows from [34, Lemma 2.2.2] and (7.37) that
$\|\max_{j}|M_{j}|\|_{\psi_{1}}\leq C\log(1+n)\max_{j}\|M_{j}\|_{\psi_{1}}\leq C\log(1+n)\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\right)|\mathscr{G}|.$ (7.40)
In conclusion, we have
@@
Conditional on $\mathcal{F}^{W}$, we have
$\mathbb{P}_{\xi}\left(\bigg{|}\sum_{j=0}^{n-1}M_{j}\bigg{|}^{p}\geq t\right)$ $=\mathbb{P}_{\xi}\left(\psi_{1}\bigg{(}\bigg{|}\sum_{j=0}^{n-1}M_{j}\bigg{|}/\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\bigg{)}\leq\psi_{1}\bigg{(}t^{\frac{1}{p}}/\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\bigg{)}\right)$ (7.42)
$\leq 1/\psi_{1}\bigg{(}t^{\frac{1}{p}}/\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\bigg{)}\leq 2\exp\left\{-\frac{t^{\frac{1}{p}}}{C(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|)|\mathscr{G}|^{\frac{1}{2}}}\right\}.$
Integrating the above equation with respect to $t$ from 00 to +$\infty$+ ∞, we have
$\mathbb{E}\left[\left|\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{j}))\mathrm{d}s-\int_{t_{i}}^{t_{i+1}}\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\mathrm{d}s\right|^{p}\bigg{|}\mathcal{F}^{W}\right]\leq C_{p}(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p})|\mathscr{G}|^{\frac{p}{2}}.$ (7.43)
Combining (7.32), (7.34), (7.35) and (7.43), using Lemma 7.4 to bound the time difference of $X^{\mathscr{G}}$ and $\widetilde{X}$ and the moments of $\sup_{t}|\widetilde{X}_{t}|$, we have
$\mathbb{E}\left[\sup_{0\leq s\leq t}\left|\Delta X_{s}\right|^{p}\right]$ $\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\mathbb{E}\left[\sup_{0\leq u\leq s}\left|\Delta X_{u}\right|^{p}\right]+\mathbb{E}\left[|X^{\mathscr{G}}_{s}-X^{\mathscr{G}}_{\delta(s)}|^{p}\right]\mathrm{d}s$ (7.44)
$\quad\quad+C_{p}|\mathscr{G}|^{\frac{p}{2}}\int_{0}^{t}\mathbb{E}[(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p})]\mathrm{d}s$
$\leq C_{p}\left|\mathscr{G}\right|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\mathbb{E}\left[\sup_{0\leq u\leq s}|\Delta X_{u}|\right]\mathrm{d}s.$
It follows from the Grönwall’s inequality that
$\mathbb{E}\left[\sup_{0\leq s\leq T}\left|\Delta X_{s}\right|^{p}\right]\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}.$ (7.45)
Now we prove the second part of this theorem. If $|\mathscr{G}|=\frac{T}{n}$, then for $p\geq 2$, $t\geq 0$,
@@
Let $t=n^{-\frac{1}{2}+\varepsilon}$,
$\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>n^{-\frac{1}{2}+\varepsilon}\right)\leq C_{p}n^{-\varepsilon p}.$ (7.47)
@@
$\sum_{n=1}^{\infty}\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>n^{-\frac{1}{2}+\varepsilon}\right)\leq C_{p}\sum_{n=1}^{\infty}n^{-1-2\varepsilon}<+\infty.$ (7.48)
@@
$J^{\mathscr{G}}-\tilde{J}=\left(\mathbb{E}\left[h(X^{\mathscr{G}}_{T})\right]-\mathbb{E}\left[h(\widetilde{X}_{T})\right]\right)+\sum_{i=0}^{n-1}\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}(r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s}))\mathrm{d}s\right]$ (7.49)
@@
As for the second term on the right hand side of (7.49), note that if $t_{i}\leq s<t_{i+1}$, then
@@
For the second part (2), we start at (7.49). By Theorem 4.2, there is a constant $C$, such that
@@
As for the second term on the right hand side of (7.49), we still have (7.51). It follows from Lemma 7.1 that
@@
Note that $\tilde{r}\in\mathcal{H}^{(l)}_{T}$, by the results of [23, Lemma 14.1.6], we have
$\left|\mathbb{E}\left[\tilde{r}(t_{i},\widetilde{X}_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]\right|\leq C\|\tilde{r}\|_{T}^{(l)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.58)
Combining these inequalities, we prove the desired result. ∎
@@
For the variance part, note that
@@
where $\Delta S_{t}:=S(t,X^{\mathscr{G}}_{t})-S(\delta(t),X^{\mathscr{G}}_{\delta(t)})$. We have
@@
$\mathrm{Var}\left(S\Box V\right)$
$\leq\mathbb{E}\left[(S\Box V)^{2}\right]$
@@
It converges to $\mathbb{E}\left[\int_{0}^{T}\widetilde{Sg}_{2}(t,\widetilde{X}_{t})\mathrm{d}t\right]$ with error bounds $C_{p}(\|S\|_{C_{p}^{2,0}}\|g_{2}\|_{C_{p}^{0,0}}+\|Sg_{2}\|_{C_{p}^{2,0}}+\|\widetilde{Sg_{2}}\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|_{C_{p}^{4}})|\mathscr{G}|$ and $C(\|S\|_{T}^{(l)}\|g_{2}\|_{T}^{(0)}+\|Sg_{2}\|_{T}^{(l)}+\|\widetilde{Sg_{2}}\|_{T}^{(l)}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|^{(l)})|\mathscr{G}|^{\mathcal{X}(l)}$ under different types of conditions by the results of Proposition 5.1. In particular, if $S$ do not depend on $a$, then
@@
Under the conditions of this proposition, we can exchange the order of integration and differentiation, therefore,
@@
Appendix A Proofs of Lemmas 3.1 and 7.2

### v2

@@ 6. Proofs of Main Results
$\leq C_{p}\mathbb{E}\left[\left(\int_{0}^{t}|\Delta X_{u}|^{2}\mathrm{d}u\right)^{\frac{p}{2}}\right]\leq C_{p}\mathbb{E}\left[\int_{0}^{t}|\Delta X_{u}|^{p}\mathrm{d}u\right],$
where $C_{p}>0$ is a generic constant that depends on $p$ and the coefficients, but is independent of the grid $\mathscr{G}$. This along with (6.26) and the regularity of $b$ and $\phi$ in (H.6) implies that
@@
$+C_{p}\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}\left(b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\right)\mathrm{d}s\right|^{p}\right].$
We now bound the last term in (6.27). For all $s\in[0,T]$, write $U_{s}:=b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})$. By (H.6), we have
@@
Denote $\mathbb{E}^{\xi}[\cdot]\coloneqq\mathbb{E}\left[\cdot|\mathcal{F}^{W}\right]$ and note that $\xi$ is independent of $\mathcal{F}^{W}$.
$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\leq C_{p}\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{\delta(u)}U_{s}\mathrm{d}s\right|^{p}\right]+C_{p}\mathbb{E}\left[\mathbb{E}^{\xi}\left[\sup_{0\leq u\leq t}\left|\int_{\delta(u)}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\right].$ (6.29)
By (6.28) and Lemma 6.4, the second term on the right hand side of (6.29) satisfies
$\mathbb{E}\left[\mathbb{E}^{\xi}\left[\sup_{0\leq u\leq t}\left|\int_{\delta(u)}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\right]$ $\leq C_{p}\mathbb{E}\left[\left(1+\sup_{0\leq t\leq T}\left|\widetilde{X}_{t}\right|^{p}+\mathbb{E}\left[\max_{1\leq j\leq n}\left(d_{A}(\phi(0,0,\xi_{j}),a_{0})\right)^{p}\right]\right)|\mathscr{G}|^{p}\right]$ (6.30)
@@
where the last inequality used the condition $|\mathscr{G}|^{1/2}\log(1+n)\leq C_{G}$.
To bound the first term on the right hand side of (6.29), for each $i=0,\ldots n-1$, let $M_{i}:=\int_{t_{i}}^{t_{i+1}}U_{s}\mathrm{d}s$ and $\mathcal{F}_{i}:=\sigma(\xi_{j}:0\leq j\leq i-1)\lor\mathcal{F}^{W}$. Since $\left\{(\sum_{j=0}^{i-1}M_{j},\mathcal{F}_{i})\right\}_{i=1}^{n}$ is a martingale, by Doob’s inequality,
$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{\delta(u)}U_{s}\mathrm{d}s\right|^{p}\right]=\mathbb{E}\left[\max_{1\leq i\leq n}\Big|\sum_{j=0}^{i-1}M_{j}\Big|^{p}\right]\leq C_{p}\mathbb{E}\left[\left|\sum_{j=0}^{n-1}M_{j}\right|^{p}\right].$ (6.31)
In the sequel, we estimate (6.31) by quantifying the tail behavior of $\left|\sum_{j=0}^{n-1}M_{j}\right|$ conditioning on the realization of the Brownian motion $W$.
Let $\|\cdot\|_{\psi_{1}}$ be the Orlicz norm on $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$, conditioning on $\mathcal{F}^{W}$. Since for all $j=0,\ldots,i-1$,
$\mathbb{E}^{\xi}[M_{j}]=\int_{t_{j}}^{t_{j+1}}\mathbb{E}^{\xi}[U_{s}]\mathrm{d}s=\int_{t_{j}}^{t_{j+1}}\mathbb{E}^{\xi}[b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})]\mathrm{d}s=0,$
and hence by [28, Theorem 6.21],
$\Big\|\sum_{j=0}^{n-1}M_{j}\Big\|_{\psi_{1}}\leq C\left(\mathbb{E}^{\xi}\left[\Big|\sum_{j=0}^{n-1}M_{j}\Big|\right]+\|\max_{j}|M_{j}|\|_{\psi_{1}}\right).$
Note that conditioning on $\mathcal{F}^{W}$, $\xi_{i}$ is the only randomness in $M_{i}$, and hence by the independent of $\xi$ and $\xi_{j}$ for $i\not=j$, $\mathbb{E}^{\xi}\left[M_{i}^{\top}M_{j}\right]=\mathbb{E}^{\xi}[M_{i}]^{\top}\mathbb{E}^{\xi}[M_{j}]=0$ for all $i\neq j$. Then using (6.28) and the square integrability of $\phi(0,0,\xi)$,
$\mathbb{E}^{\xi}\bigg[\Big|\sum_{j=0}^{n-1}M_{j}\Big|\bigg]\leq\sqrt{\mathbb{E}^{\xi}\bigg[\Big|\sum_{j=0}^{n-1}M_{j}\Big|^{2}\bigg]}=\sqrt{\sum_{j=0}^{n-1}\mathbb{E}^{\xi}\left[|M_{j}|^{2}\right]}\leq C\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\right)|\mathscr{G}|^{\frac{1}{2}}.$
By (6.28) and (H.6),
$\|M_{j}\|_{\psi_{1}}\leq C\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\|1\|_{\psi_{1}}+\|d_{A}(\phi(0,0,\xi),a_{0})\|_{\psi_{1}}\right)|\mathscr{G}|\leq C\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\right)|\mathscr{G}|,$ (6.32)
which along with Lemma 6.4 implies that $\left\|\max_{j}|M_{j}|\right\|_{\psi_{1}}\leq C\log(1+n)\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\right)|\mathscr{G}|$, and
@@
For all $t>0$, by the conditional Markov inequality and $\mathbb{E}^{\xi}[\psi_{1}(|\sum_{j=1}^{n-1}M_{j}|/\|\sum_{j=1}^{n-1}M_{j}\|_{\psi_{1}})]\leq 1$,
$\mathbb{P}^{\xi}\left(\bigg|\sum_{j=0}^{n-1}M_{j}\bigg|^{p}\geq t\right)$ $=\mathbb{P}^{\xi}\left(\psi_{1}\left(\frac{\left|\sum_{j=0}^{n-1}M_{j}\right|}{\left\|\sum_{j=0}^{n-1}M_{j}\right\|_{\psi_{1}}}\right)+1\leq\psi_{1}\left(\frac{t^{\frac{1}{p}}}{\left\|\sum_{j=0}^{n-1}M_{j}\right\|_{\psi_{1}}}\right)+1\right)$
$\leq 2\left(\psi_{1}\left(\frac{t^{\frac{1}{p}}}{\left\|\sum_{j=0}^{n-1}M_{j}\right\|_{\psi_{1}}}\right)+1\right)^{-1}\leq 2\exp\left\{-\frac{t^{\frac{1}{p}}}{C(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|)|\mathscr{G}|^{\frac{1}{2}}}\right\},$
where the last inequality used that $(\psi_{1}(x)+1)^{-1}=\exp(-x)$ for all $x>0$, and (6.33). Thus
$\mathbb{E}^{\xi}\left[\bigg|{\sum_{j=0}^{n-1}M_{j}}\bigg|^{p}\right]=\int_{0}^{\infty}\mathbb{P}^{\xi}\left(\bigg|{\sum_{j=0}^{n-1}M_{j}}\bigg|^{p}\geq t\right)\mathrm{d}t\leq C_{p}\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p}\right)|\mathscr{G}|^{\frac{p}{2}}.$ (6.34)
Combining (6.27), (6.29), (6.30), (6.31) and (6.34), using Lemma 6.5 to bound the time difference of $X^{\mathscr{G}}$ and $\widetilde{X}$ and the moments of $\sup_{t}|\widetilde{X}_{t}|$, we have
$\mathbb{E}\left[\sup_{0\leq s\leq t}\left|\Delta X_{s}\right|^{p}\right]$ $\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\left(\mathbb{E}\left[\sup_{0\leq u\leq s}\left|\Delta X_{u}\right|^{p}\right]+\mathbb{E}\left[|X^{\mathscr{G}}_{s}-X^{\mathscr{G}}_{\delta(s)}|^{p}\right]\right)\mathrm{d}s$ (6.35)
$\quad+C_{p}|\mathscr{G}|^{\frac{p}{2}}\int_{0}^{t}\mathbb{E}\left[1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p}\right]\mathrm{d}s$
$\leq C_{p}\left|\mathscr{G}\right|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\mathbb{E}\left[\sup_{0\leq u\leq s}|\Delta X_{u}|^{p}\right]\mathrm{d}s,$
which yields the desired estimate $\mathbb{E}\left[\sup_{0\leq s\leq T}\left|\Delta X_{s}\right|^{p}\right]\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}$ due to Grönwall’s inequality.
It remains to prove the almost sure convergence with $|\mathscr{G}|=\frac{T}{n}$. For $p\geq 2$ and $t\geq 0$,
@@
Given $\varepsilon\in(0,1)$, setting $t=n^{-\frac{1}{2}+\varepsilon}$ in (6.36) yields for all $p\geq 2$,
$\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>n^{-\frac{1}{2}+\varepsilon}\right)\leq C_{p}n^{-\varepsilon p}.$
@@
$\sum_{n=1}^{\infty}\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>n^{-\frac{1}{2}+\varepsilon}\right)\leq C_{p}\sum_{n=1}^{\infty}n^{-1-2\varepsilon}<+\infty.$
@@
6.2. Proof of Corollary 4.1
Proof of Corollary 4.1.
Recall that $\mathbb{P}^{W}$ is the conditional probability measure given $\mathcal{F}^{\xi}$. Under the measure $\mathbb{P}^{W}$, $\{f(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi))\}_{k=1}^{m}$ are independent random variables taking values in $\|f\|_{\infty},\|f\|_{\infty}]$, where $\|f\|_{\infty}$ is the sup-norm of $f$. By Hoeffding’s inequality, for all $t\geq 0$ and $m,n\in{\mathbb{N}}$,
$\mathbb{P}^{W}\left(\left|\frac{1}{m}\sum_{k=1}^{m}\left(f\left(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi)\right)-\mathbb{E}^{W}\left[f\left(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi)\right)\right]\right)\right|\geq t\right)\leq 2\exp\left(-\frac{2t^{2}m}{4\|f\|_{\infty}^{2}}\right),$ (6.37)
For a given $\varrho\in(0,1)$, choosing $t$ such that $\varrho=2\exp(-t^{2}m/(2\|f\|_{\infty}^{2}))$ yields
$\mathbb{P}^{W}\left(\left|\frac{1}{m}\sum_{i=1}^{m}f\left(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi)\right)-\mathbb{E}^{W}\left[f\left(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi)\right)\right]\right|\geq\sqrt{2}\|f\|_{\infty}\frac{1}{\sqrt{m}}\ln(2/\varrho)\right)\leq\varrho.$ (6.38)
This along with Theorem 4.3 shows that there exists a constant $C\geq 0$ such that for all $n,m\in{\mathbb{N}}$ and $\varrho\in(0,1)$,
$\mathbb{P}\left(\bigg|\hat{I}_{m,n}-\mathbb{E}[f(\widetilde{X}_{T})]\bigg|\geq C\left(1+\sqrt{\ln(2/\varrho)}\right)(n^{-1/2}+m^{-1/2})\right)\leq\varrho.$ (6.39)
The desired result follows directly by choosing $n$ and $m$ as in the statement. ∎
@@
$J^{\mathscr{G}}-\tilde{J}=\left(\mathbb{E}\left[h(X^{\mathscr{G}}_{T})\right]-\mathbb{E}\left[h(\widetilde{X}_{T})\right]\right)+\sum_{i=0}^{n-1}\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}(r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s}))\mathrm{d}s\right].$ (6.40)
@@
For the second term on the right hand side of (6.40), note that if $t_{i}\leq s<t_{i+1}$, then
@@
For the second part (2), we start at (6.40). By Theorem 4.2, there is a constant $C\geq 0$ such that
@@
For the second term on the right hand side of (6.40), we still have (6.42). By Lemma 6.2,
@@
Note that $\tilde{r}\in\mathcal{H}^{(l)}_{T}$, by [25, Lemma 14.1.6], $\left|\mathbb{E}\left[\tilde{r}(t_{i},\widetilde{X}_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]\right|\leq C\|\tilde{r}\|_{T}^{(l)}|\mathscr{G}|^{\mathcal{X}(l)}.$ Combining these inequalities yield the desired result. ∎
@@
For the variance, note that
@@
[…]
@@
$\mathrm{Var}\left(S\Box V\right)\leq\mathbb{E}\left[(S\Box V)^{2}\right]$
@@
It converges to $\mathbb{E}\left[\int_{0}^{T}\widetilde{Sg}_{2}(t,\widetilde{X}_{t})\mathrm{d}t\right]$ with error bounds $C_{p}(\|S\|_{C_{p}^{2,0}}\|g_{2}\|_{C_{p}^{0,0}}+\|Sg_{2}\|_{C_{p}^{2,0}}+\|\widetilde{Sg_{2}}\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|_{C_{p}^{4}})|\mathscr{G}|$ and $C(\|S\|_{T}^{(l)}\|g_{2}\|_{T}^{(0)}+\|Sg_{2}\|_{T}^{(l)}+\|\widetilde{Sg_{2}}\|_{T}^{(l)}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|^{(l)})|\mathscr{G}|^{\mathcal{X}(l)}$ under different types of conditions by the results of Proposition 5.1. In particular, if $S$ does not depend on $a$, then
@@
Under the given conditions, we can exchange the order of integration and differentiation, and obtain
@@
Acknowledgments
Yanwei Jia is supported by the Start-up Fund at The Chinese University of Hong Kong, and Hong Kong Research Grants Council (RGC) - Early Career Scheme (ECS) 24211124. We are especially indebted to the two anonymous referees for their constructive and detailed comments that have led to an improved version of the paper. We thank conference participants at IASM Workshop on “Bridging Theory and Practice in Finance: Stochastic Analysis and Machine Learning”, Workshop on Stochastic Control and Reinforcement Learning at Harbin Institute of Technology, 2nd ETH-Hong Kong-Imperial Mathematical Finance Workshop, and 15th POMS-HK International Conference for their helpful discussions.
Appendix A Proofs of Lemmas 3.1, 6.1 and 6.3